A)What is the total heat capacity of the calorimet

A)What is the total heat capacity of the calorimeter?

B)A 1.450 ?g sample of a new organic substance is combusted in the same calorimeter. The temperature of the calorimeter increases from 22.14 ?C to 27.09 ?C. What is the heat of combustion per gram of the new substance?
Answer #1

Given that

Mass of benzoic acid = 2.780 g

Heat of combustion of benzoic acid= 26.38 KJ/g

Therefore

(1) Heat released = mass of benzoic acid x heat of combustion

= 2.780 x 26.38

= 73.3364 kJ


Heat capacity of calorimeter = heat released/temperature change of calorimeter

= 73.3364 /(29.98 - 21.40)

= 73.3364 /(8.58)

= 8.547 kJ/C

(2) Heat released = heat capacity x temperature change of calorimeter

= 8.547 x (27.09 - 22.14) = 41.7493 kJ


Heat of combustion = heat released/mass of substance

= 41.7493/1.450

= 28.8 kJ/g